2x+3=4x^2: 2x+3=4x^2
Answer by ankor@dixie-net.com(4508)

2x+3=4x^2
:
You can rearrange this is the quadratic form:
4x^2 - 2x - 3 = 0
:
Use the quadratic formula:
x = (-b +- sqrt( b^2-4*a*c ))/(2*a)
In your problem: a=4; b=-2; c=-3
x = ((-2) +- sqrt( -2^2-4*4*-3 ))/(2*4)
:
x = (2 +- sqrt(4 - (-48) ))/(8)
x = (2 +- sqrt(4 + 48))/(8)
x = (2 +- sqrt(52 ))/(8)
Two solutions:
x = (2 + 7.211)/8
x = (9.211)/8
x = 1.151
and
x = (2 - 7.211)/8
x = (-5.211)/8
x = -.651